### v1

@@ 4.4.3 Selection of top_k Functions
[…]
@@ 4.6 Performance Metric
In our application scenarios (XSS/SQLi detection), we give more weight to false negatives (resulting in low recall) with respect to false positives (resulting in low precision), since a non-detected attack can cause much more damage than a benign request detected as an attack. For this reason, the main performance metric used in our empirical study is the F2-Score, referred to as F2 for simplicity ($\mathcal{M}=F_{2}$), which gives double importance to recall than to precision: $F_{2}=5p*r/(4p+r)$, with $p$ and $r$ indicating precision and recall, respectively. To mitigate a possible threat to the construct validity associated with the choice of the performance metric F2, which is strongly related to the specific domain of the study, we replicated all experiments using accuracy and F1-Score, referred to as F1 and defined as $F_{1}=2p*r/(p+r)$, as alternative performance metrics. Accuracy is a valuable choice since the datasets used in this study are well-balanced. However, since in a real-world scenario positive samples are much fewer than the negative ones, we replicated all the experiments also using F1-Score, a metric that ignores the false negatives and hence, differently from accuracy, is not sensitive to the balance of positive vs negative data. Results are consistent and largely independent of the choice of the performance metric. The interested reader can find the additional results obtained with $\mathcal{M}=Accuracy$ and $\mathcal{M}=F1$ in the replication package (see Section Data Availability).
@@ 5.1 RQ1 (RAG Benefit)
Figure 3 illustrates the impact of RAG on function generation for XSS and SQLi in the NTD scenario, showing the F2 differences between the configurations with and without RAG, across all possible configurations. Each bar represents the average F2 score difference for a specific Model-Temperature pair when using RAG, compared to the same pair without RAG. Results indicate that employing RAG generally enhances the performance of function generation for both tasks. The number of Model-Temperature pairs benefiting from RAG is much larger than the number of pairs showing degradation, and the improvements are statistically significant, as evidenced by $p$-values ($\approx 10^{-66}$ for XSS and $\approx 10^{-24}$ for SQLi) below the standard threshold of 0.05.
@@
Turning to the TDA scenario where we can select the best Generated Function Run using $val\_set$, we observe that $U^{best}$ takes always advantage of RAG as well as Few-shot examples: for XSS, $U^{best}=$ (GPT-4T, 0.0, 10, T), and for SQLi, $U^{best}=$ (GPT-4T, 0.0, 2, T). This indicates that while the use of few-shot examples may not consistently enhance performance in general (as seen in the NTD scenario), the best configuration selected with $val\_set$ (in the TDA scenario) leverages the benefits of using both RAG and Few-shot examples.
@@ 5.2 RQ2 (Impact of Self-Ranking)
Let us consider the NTD scenario. The violin plots shown in Figure 5 depict the effect of Self-Ranking, i.e., $top\_k$ selection across three values of $k$ for the two tasks, considering all the possible pairs of function/synthetic dataset configurations. Overall, we observe a clear improvement, particularly for XSS. The region between quartiles for XSS falls largely between a 20%pt and 40%pt improvement, with an average improvement of 37%pt and an improvement that affects 98% of the cases. While the improvement for SQLi is less pronounced, we still observe improvements in 73% cases, with an average improvement of 6%pt. Additionally, we can see that, as $k$ increases, the average improvement decreases for both tasks, but the gains become more stable. The improvements given by the usage of $top\_k$ selection are statistically significant: the $p$-values obtained with the Wilcoxon signed-rank test are below the threshold of 0.05 for both tasks and all the values of $k$ [^10] Specifically, $p$-value is 00 for all the $k$s for XSS. For SQLi, it was 00 when $k=1$ and $k=2$ and $\approx 10^{-247}$ when $k=5$..
@@
Answer to RQ2: Utilizing the Self-Ranking mechanism that selects $top\_k$ functions leads to statistically significant improvements on both tasks.
@@
Baseline 0.630 Baseline 0.800
@@ 5.3 RQ3 (Comparison with SOTA)
[…]
Table 4 presents the results of the comparison. We observe a significant improvement of 34%pt and 18%pt for XSS and SQLi respectively, when compared to the baseline. There is a slight decrease in F2 compared to SOTA models, with a 3.15%pt and 0.83%pt drop for XSS and SQLi, respectively. We argue that the slight performance gap between our approach and SOTA models is understandable, given our approach’s training-free nature and direct applicability to multiple tasks. Moreover, these results provide empirical support for our claim that incorporating external knowledge and Self-Ranking is essential for LLMs to achieve competitive performance with SOTA models.
@@ 5.4 RQ4 (Transferability)
Table 6 presents the transferability results. The ‘$F2(U^{best},S^{best})$’ columns indicate F2 computed on the original task with its best configuration, ‘Avg. F2’ columns represent the average F2 computed across all the $U-S$ pairs for a given $k$, and the ‘$F2(U^{transf},S^{transf})$’ columns show F2 computed using transferred configurations. Comparing the transferred configuration’s performance to the average column provides a good estimate of the benefits of configuration transfer over a mere random selection of a configuration for a new, unseen task.
@@ 6 Discussion
With RQs 1 and 2, we highlighted the benefits of using RAG and Self-Ranking, highlighting the necessity to incorporate external knowledge about the attacks to generate robust detectors, and exploiting the multiple reasoning-paths of the LLM to select the $K$ top-ranked detectors. However, these results do not offer explicit insights into the best choices for the other parameters, such as LLM or temperature. While this question may have a trivial answer in the TDA scenario where developers can experiment and evaluate different configurations, the NTD scenario presents a more challenging situation. On one hand, our transferability study (RQ4) demonstrated that the best-performing combinations for one task exhibit strong performance on the other task. Therefore, the combinations from another task could serve as a reasonable starting point for selecting the inner parameters when approaching a new task.
When transferability is not possible (e.g., because developers cannot access the LLM identified via transferability), developers have still to choose a model and a temperature. We can support their choice by conditioning the results of our experiments on model or on temperature. Results conditioned for each model are shown in Figure 6, which presents the average F2 achieved by different LLMs across the two tasks. These results indicate the strong performance of GPT-4T and GPT-4, which is unsurprising given their inclusion within the $U^{best}$ in the earlier experiments. On the other hand, Mixtral consistently underperforms relative to other LLMs in both tasks. While these plots exhibit some correlation with the HumanEval scores reported in Table 1, it is important to note that models like Haiku, which achieved decent HumanEval scores, were discarded due to their inability to successfully complete the task (see Section 4.3). This observation aligns with our hypothesis that while HumanEval effectively assesses the reasoning capabilities of LLMs, it may not directly translate to the depth of knowledge required for more specialized tasks, such as secure code generation.
@@ 7 Related Work
SVEN he2023large is a learning-based approach that aims to guide LLM’s code generation towards satisfying a given property, exploiting Prefix-Tuning li2021prefix to fine-tune the LLM. While SVEN demonstrated promising results in preventing LLMs from introducing vulnerabilities, it requires a fine-tuning procedure based on a curated training set, hindering its efficiency and applicability to closed LLMs, hence limiting its scope.
@@ 8 Threats to Validity
Internal validity. LLMs are non-deterministic, hence we repeated our function and synthetic dataset generation experiments 40 and 10 times respectively (the disparity is due to the higher stability of the latter experiments). Indeed, we adopt Self-Ranking to exploit the non-deterministic nature of LLMs. While different LLM configurations and selection of RAG sources may yield varying results, our choices were based on documentation and existing best practices. Data leakage from LLM training could be a concern, but given the complexity of the task and the presence of only a few generated functions with perfect performance, we believe it was not an important factor in our experiment.
@@ 9 Conclusion
The authors declare that they do not have any known relationship or competing interests that could have influenced this paper. The authors declare that their research for the current work did not involve Human Participants or Animals.

### v2

@@ 4.4.3 Selection of top_k Functions
We select the top_k functions of a Generated Functions Run $U$ using a dataset $d$ and a performance metric $\mathcal{M}$ (e.g., F2-Score) by sorting the functions in $U$ by decreasing $\mathcal{M}$ and including only the first $k$ functions in the sorted list: $\textit{top\_k}(U,d,\mathcal{M})=\textit{sort}(U,\mathcal{M}(\cdot,d))[1\mathpunct{:}k]$.
Note that in the event of a tie for the $k^{th}$ position, one element is randomly chosen. In the TDA scenario, we can set $d=\textit{val\_set}$ to select the top_k functions. In the NTD scenario, we use a Synthetic Dataset Run as $d$, exploiting Self-Ranking. In Figure 2, the TDA scenario is shown on the top-right, with $d$ the Validation Dataset. The NTD scenario is shown on the bottom-right, with $d$ equal to each of the generated datasets $S_{1}$ and $S_{2}$.
The usage of Self-Ranking requires black-box access to the LLM for Dataset Generation, as it is purely based on prompting, while top_k selection is performed without any query to the LLM, as the ranking score is based on the generated function’s performance on the synthetic dataset.
@@ 4.6 Performance Metric
In our application scenarios (XSS/SQLi detection), we give more weight to false negatives (resulting in low recall) with respect to false positives (resulting in low precision), since a non-detected attack can cause much more damage than a benign request detected as an attack. For this reason, the main performance metric used in our empirical study is the F2-Score, referred to as F2 for simplicity ($\mathcal{M}=F_{2}$), which gives double importance to recall than to precision: $F_{2}=5p*r/(4p+r)$, with $p$ and $r$ indicating precision and recall, respectively. The choice of a metric that gives more weight to the recall than to the precision is related to the fact that, in many domains, an undetected attack is considered more dangerous than a false positive. To mitigate a possible threat to the construct validity associated with the choice of the performance metric F2, which is strongly related to the specific domain of the study, we replicated all experiments using accuracy and F1-Score, referred to as F1 and defined as $F_{1}=2p*r/(p+r)$, as alternative performance metrics. Accuracy is a valuable choice since the datasets used in this study are well-balanced. However, since in a real-world scenario positive samples are much fewer than the negative ones, we replicated all the experiments also using F1-Score, a metric that ignores the false negatives and hence, differently from accuracy, is not sensitive to the balance of positive vs negative data. Results are consistent and largely independent of the choice of the performance metric. The interested reader can find the additional results obtained with $\mathcal{M}=Accuracy$ and $\mathcal{M}=F1$ in the replication package.
@@ 5.1 RQ1 (RAG Benefit)
Figure 3 illustrates the impact of RAG on function generation for XSS and SQLi in the NTD scenario, showing the F2 differences between the configurations with and without RAG, across all possible configurations. Each bar represents the average F2 score difference for a specific Model-Temperature pair when using RAG, compared to the same pair without RAG. Results indicate that employing RAG generally enhances the performance of function generation for both tasks. The number of Model-Temperature pairs benefiting from RAG is much larger than the number of pairs showing degradation, and the improvements are statistically significant, as evidenced by $p$-values ($\approx 10^{-66}$ for XSS and $\approx 10^{-24}$ for SQLi) below the standard threshold of 0.05. Furthermore, models such as GPT-4, which demonstrated superior performance without RAG, also exhibit greater improvements when RAG is employed. This suggests that the baseline knowledge of these models is not only more robust but their capacity to utilize the contextual information provided by RAG is also more effective.
(a) XSS Detection
(b) SQLi Detection
@@
(a) XSS Detection
(b) SQLi Detection
Figure 5: Difference between the F2 of Generated Function Runs with Few-shot Prompt and the F2 of Generated Function Runs with Zero-shot Prompt, for XSS detection (left) and SQLi detection (right).
However, as illustrated in Figure 5, Few-shot examples offer a significant advantage when RAG is not employed. This is likely due to the well-documented challenge LLMs face in effectively retrieving and utilizing specific knowledge from large contexts liu-etal-2024-lost . In the absence of RAG, Few-shot examples play a crucial role in guiding the model. However, when RAG is used, the extensive amount of provided knowledge appears to overshadow the influence of the Few-shot examples, rendering them less effective within the broader context.
Turning to the TDA scenario where we can select the best Generated Function Run using val_setval\_set, we observe that $U^{best}$ takes always advantage of RAG as well as Few-shot examples: for XSS, $U^{best}=$ (GPT-4T, 0.0, 10, T), and for SQLi, $U^{best}=$ (GPT-4T, 0.0, 2, T). For brevity, we omit the full results table for the TDA scenario here, but it is included in the replication package. This indicates that while the use of few-shot examples may not consistently enhance performance in general (as seen in the NTD scenario), the best configuration selected with val_setval\_set (in the TDA scenario) leverages the benefits of using both RAG and Few-shot examples.
@@ 5.2 RQ2 (Impact of Self-Ranking)
Let us consider the NTD scenario. The violin plots shown in Figure 6 depict the effect of Self-Ranking, i.e., top_ktop\_k selection across three values of $k$ for the two tasks, considering all the possible pairs of function/synthetic dataset configurations. Overall, we observe a clear improvement, particularly for XSS. The region between quartiles for XSS falls largely between a 20%pt and 40%pt improvement, with an average improvement of 37%pt (up to 71%pt) and an improvement that affects 98% of the cases. While the improvement for SQLi is less pronounced, we still observe improvements in 73% cases, with an average improvement of 6%pt (up to 43%pt). Additionally, we can see that, as $k$ increases, the average improvement decreases for both tasks, but the gains become more stable. The improvements given by the usage of top_ktop\_k selection are statistically significant: the $p$-values obtained with the Wilcoxon signed-rank test are below the threshold of 0.05 for both tasks and all the values of kk [^14] Specifically, $p$-value is 0 for all the $k$s for XSS. For SQLi, it was 0 when $k=1$ and $k=2$ and $\approx 10^{-247}$ when $k=5$..
@@
The Synthetic Datasets generated with the LLMs proved to be effective in improving the final performance, exploiting Self-Ranking. However, for completeness, it is also interesting to directly assess the quality of the generated Synthetic Datasets. To have a better overview of the quality of those datasets, we randomly sampled 100 Malicious and Benign samples from various Synthetic Datasets and we asked three human evaluators with many-year expertise in the security domain to label them as Malicious or Benign.
Given an example of Malicious/Benign HTTP/SQLi payload generated inside a Synthetic Dataset, the task for the evaluators was to label it as Malicious or Benign in a simple multiple-choice question. The question was preceded by a description of the task and a few ground-truth examples (both for XSS and SQLi). An example of a Benign HTTP Payload is http://www.github.com/user/repo; http://example.com/?search=alert(’XSS’) is an example of a Malicious HTTP payload, while UPDATE users SET password = ’1’ or ’1’ = ’1’, is_admin = 1 and SELECT * FROM users WHERE id = 1 are two examples of, respectively, a Malicious and a Benign SQL query.
After selecting the annotation as the majority vote among the three annotators, the agreement with the label assigned by the LLM while generating the samples was 98%. This result underlines that the generated Synthetic Datasets are not just a good tool to be exploited by Self-Ranking, but they are a good-quality artifact that can be potentially exploited in a larger set of tasks.
Answer to RQ2: Utilizing the Self-Ranking mechanism that selects top_ktop\_k functions leads to statistically significant improvements on both tasks.
@@
Baseline (Few-shot) 0.922 Baseline (Few-shot) 0.882
Baseline (Basic Prompt) 0.630 Baseline (Basic Prompt) 0.800
@@ 5.3 RQ3 (Comparison with SOTA)
To compare ours with learning-based SOTA techniques, we select $U^{best}$ and top_ktop\_k functions based on val_setval\_set. For XSS, $U^{best}$ is $(\textsc{GPT-4T},0.0,10,T)$, while for SQLi, it is $(\textsc{GPT-4T},0.0,2,T)$. Due to lack of space, we omit the table with all the results, which is anyway available in the replication package. We establish two baselines to better understand the improvement offered by adopting our approach. The first baseline is obtained using Basic Prompt, without Few-shot examples and RAG. It represents the expected results that can be obtained by generating detectors without exploiting any of the techniques presented in this work. The second one is obtained by considering also the usage of Few-shot examples, and it represents the expected results that can be obtained without exploiting any of the techniques that are the direct contributions of this work.
Table 4 presents the results of the comparison. We observe a significant improvement of 34%pt and 18%pt for XSS and SQLi respectively, when compared to the Basic Prompt baseline. The improvement with respect to the Few-shot Baseline is of 4.3%pt and 10%pt for XSS and SQLi respectively. There is a slight decrease in F2 compared to SOTA models, with a 3.15%pt and 0.83%pt drop for XSS and SQLi, respectively. We argue that the slight performance gap between our approach and SOTA models is understandable, given our approach’s training-free nature and direct applicability to multiple tasks. Moreover, these results provide empirical support for our claim that incorporating external knowledge and Self-Ranking is essential for LLMs to achieve competitive performance with SOTA models.
@@ 5.4 RQ4 (Transferability)
Table 6 presents the transferability results. The $F2(U^{best},S^{best})$’ columns indicate F2 computed on the original task with its best configuration, ‘Avg. F2’ columns represent the average F2 computed across all the USU-S pairs for a given $k$, and the $F2(U^{transf},S^{transf})$’ columns show F2 computed using transferred configurations. Comparing the transferred configuration’s performance to the average column provides a good estimate of the benefits of configuration transfer over a mere random selection of a configuration for a new, unseen task.
@@ 6 Discussion
With RQs 1 and 2, we highlighted the benefits of using RAG and Self-Ranking, highlighting the necessity to incorporate external knowledge about the attacks to generate robust detectors, and exploiting the multiple reasoning-paths of the LLM to select the $K$ top-ranked detectors. However, these results do not offer explicit insights into the best choices for the other parameters, such as LLM or temperature. While this question may have a trivial answer in the TDA scenario where developers can experiment and evaluate different configurations, the NTD scenario presents a more challenging situation. On one hand, our transferability study (RQ4) demonstrated that the best-performing combinations for one task exhibit strong performance on the other task. Therefore, the combinations from another task could serve as a reasonable starting point for selecting the inner parameters when approaching a new task. Since we considered just two tasks, results are not enough to show generalizability to tasks different from the considered ones. Still, they suggest potential transferability under similar task characteristics. When transferability is not possible (e.g., because developers cannot access the LLM identified via transferability), developers have still to choose a model and a temperature. We can support their choice by conditioning the results of our experiments on model or on temperature. Results conditioned for each model are shown in Figure 7, which presents the average F2 achieved by different LLMs across the two tasks. These results indicate the strong performance of GPT-4T and GPT-4, which is unsurprising given their inclusion within the $U^{best}$ in the earlier experiments. On the other hand, Mixtral consistently underperforms relative to other LLMs in both tasks. While these plots exhibit some correlation with the HumanEval scores reported in Table 1, it is important to note that models like Haiku, which achieved decent HumanEval scores, were discarded due to their inability to successfully complete the task (see Section 4.3). This observation aligns with our hypothesis that while HumanEval effectively assesses the reasoning capabilities of LLMs, it may not directly translate to the depth of knowledge required for more specialized tasks, such as secure code generation.
@@
Regarding the knowledge sources used in RAG, we selected them manually. However, assuming the deployment of the proposed method in a real-world setting, we can envision an automated retrieval process that leverages a predefined set of reliable sources, such as those from OWASP, to extract relevant information without the need for manual curation.
Our approach assumes that a developer aiming to use the LLM generation process can identify and access a suitable RAG source for their target task. They must also be able to retrieve or create a few representative examples to guide the generation process and either craft an effective task-specific prompt from scratch or adapt one of our provided prompts (e.g., those designed for XSS or SQLi) to suit their specific use case.
When considering the choice between our LLM-based approach and SOTA trainable models like DNN, MLP, or CNN, the decision depends on the availability of a labeled dataset. Our results demonstrate that, in the absence of a large dataset, our methodology can achieve results comparable to SOTA without this prerequisite.
@@ 7 Related Work
SVEN he2023large is a learning-based approach that aims to guide LLM’s code generation towards satisfying a given property, exploiting Prefix-Tuning li2021prefix to fine-tune the LLM. While SVEN demonstrated promising results in preventing LLMs from introducing vulnerabilities, it requires a fine-tuning procedure based on a curated training set, hindering its efficiency and applicability to closed LLMs, hence limiting its scope.To the best of our knowledge, our work is the first work that generates attack detectors for SQLi and XSS instead of training a DNN to solve these tasks or focusing on vulnerability detection in existing code.
@@ 8 Threats to Validity
Internal validity. LLMs are non-deterministic, hence we repeated our function and synthetic dataset generation experiments 40 and 10 times respectively (the disparity is due to the higher stability of the latter experiments). Indeed, we adopt Self-Ranking to exploit the non-deterministic nature of LLMs. While different LLM configurations and selection of RAG sources may yield varying results, our choices were based on documentation and existing best practices.
@@
Our approach was evaluated on two tasks where high-quality documentation–used as RAG sources in our experiments–was available. However, for other tasks, obtaining knowledge sources of comparable quality might be challenging. In such cases, the overall performance of our method may change. Specifically, the positive impact of RAG is likely to be reduced, while Self-Ranking retains its potential to contribute.
Since we are using publicly available datasets, data contamination is a possible concern in this work, as in any LLM-based work. However, since the functions generated by the LLM alone are far from perfect, it is reasonable to conjecture that data contamination is not a significant issue in our case.
@@ 9 Conclusion
The authors declare that they do not have any known relationship or competing interests that could have influenced this paper. The authors declare that the Human involved in the study for the assessment of the quality of the Generated Synthetic Dataset signed an informed consent approved by the Ethical Committee of Università della Svizzera Italiana. The authors declare that their research for the current work did not involve Animals.